# Fairness in AI-Driven Recruitment: Challenges, Metrics, Methods, and Future Directions

### v2

@@ IV Bias Detection and Mitigation
To ensure fairness in AI-driven hiring, it is crucial to implement methods for detecting and measuring bias and mitigating it at each stage of the hiring pipeline. In this section, we review the existing work in AI fairness, focusing on methods for measuring bias, strategies for mitigating bias, and the current limitations and challenges in this field.
IV-A Fairness Metrics
In AI fairness research, various definitions of fairness have been proposed to assess AI systems effectively. In this subsection, we explore core fairness metrics particularly relevant to recruitment, as defined in prior research [66, 71, 74].
@@ IV-A Fairness Metrics
    Fairness through unawareness: This metric serves as a baseline by stating that an algorithm is fair if it does not explicitly use protected attributes $A$ in its training [125]. The classifier $C$, defined as $C:X\setminus A\to Y$, operates on a set of features $X$ excluding any protected attributes $A$. Here, $Y$ represents the potential outcome classes. While this approach aims to prevent direct discrimination by ensuring the classifier is ’unaware’ of protected characteristics, it has notable shortcomings. In the hiring context, where $A$ might vary significantly between different occupations, expert knowledge is essential to accurately identify and exclude these attributes. Moreover, even when $A$ is omitted from $X$, other features in $X$ might act as proxies to $A$, leading to a classifier that is nominally fair while still harboring discriminatory information. For example, while gender may not be explicitly included, participation in a women’s coding bootcamp listed in job history could still suggest the candidate’s gender, thus acting as a proxy and influencing the hiring decision. Additionally, educational background might act as a proxy for socioeconomic status or ethnic background, especially if certain schools predominantly serve specific demographic groups.
@@
    Demographic parity: Also known as statistical parity, this metric stipulates that the acceptance rates for different demographic groups should be approximately equal. This requirement aligns with fair hiring practices mandated by laws such as the four-fifths rule (see Section II-B). Mathematically, let $C$ be a classifier predicting a binary outcome $Y$. For any two groups $a$ and $b$, the acceptance rates must be within a specified percentage $p$ of each other, where $\epsilon=\frac{p}{100}$ and $\epsilon\in[0,1]$ [71]:
@@
    This metric does not address the trade-off between false positives and false negatives directly but focuses on ensuring that all candidates who are qualified have an equal chance of being selected, regardless of group membership.
@@
    These conditions ensure that the model’s predictions do not disproportionately benefit or penalize any group, regardless of the actual outcome.
@@
    Counterfactual fairness: Motivated by the idea that decisions made by an algorithm should not change if a protected attribute $A$ were hypothetically altered, counterfactual fairness enhances transparency by demanding robustness in predictions against changes in $A$. Defined by Kusner et al. [125], a classifier $C$ achieves counterfactual fairness if, for any individual with features $X$ including a protected attribute $A$, the outcome remains the same in both the actual and a counterfactual scenario where $A$ is altered. The mathematical definition is:
@@
    Individual fairness (versus group fairness): Unlike group fairness metrics such as demographic parity, accuracy parity, and predictive rate parity which ensure fairness by equalizing statistical measures across groups, individual fairness focuses on fairness at the individual level. This metric asserts that similar individuals should receive similar outcomes. It addresses the key limitation of group fairness metrics, which ensure fairness across groups but may still permit unfairness within groups [104]. Specifically, individual fairness ensures that any two individuals $i$ and $j$ who are similar with respect to the task at hand should be treated similarly [126].
@@
    Preference-informed individual fairness: This fairness metric extends individual fairness by considering the diverse preferences of individuals, such as job location, industry, salary range, and work hours. Preference-informed individual fairness (PIIF) integrates the concept of envy-freeness, which ensures that no individual would prefer another’s outcome to their own under similar conditions [126]. Thus, PIIF demands that outcomes not only meet individual fairness criteria but also align with the individual preferences to mitigate envy, ensuring each decision made by the model is both fair and personally satisfactory to the individuals involved.
@@
    Multi-sided fairness: Modern recruitment platforms necessitate a multi-sided approach to fairness, considering the provider (i.e., employers posting job advertisements), the consumer (i.e., jobseekers), and the platform itself (i.e., the job posting recommendation or screening system) [127]. Prioritizing the interests of only one party, such as focusing solely on the employer (producer-focused) or the jobseeker (consumer-focused), can lead to unfair outcomes for the other. For instance, a consumer-focused approach, or C-fairness, might mandate that job ads display similar salary ranges to all users regardless of protected group membership. Conversely, a producer-focused approach, or P-fairness, might require that ads from minority-owned businesses be shown to qualified jobseekers at comparable rates to those from other businesses [128, 129]. This concept has been applied in various online platforms such as e-commerce and ride-hailing, where fairness is treated as a multi-objective optimization problem, balancing accuracy and fairness [127, 130, 128, 129].
@@
    Procedural fairness: Procedural fairness, one of the three types of organizational justice discussed in Section II-A—alongside distributive and interactional justice—focuses on the fairness of processes within an organization, such as recruitment. Grgic-Hlaca et al. [44] propose using procedural fairness to evaluate recruitment processes. Specifically, the fairness of a classifier $C$ and its feature set $X$ is assessed based on the proportion of users who perceive the use of certain features in the decision-making process as fair. Essentially, this approach allows users to actively evaluate the fairness of decisions. However, perceptions of fairness may vary over time and are influenced by the users’ awareness of potential disparities caused by specific features in the model [44].
@@
    Fairness in regression: While many established fairness metrics focus on classification problems where decisions are typically binary (e.g., accept or reject), regression models are also prevalent in hiring processes, such as scoring video interviews [131]. These models, however, have received less attention regarding fairness. According to Barocas et al. [70], fairness in regression is guided by three principles:
    • Independence — ensuring that a sensitive attribute $A$ and the regression output $R$ are independent, denoted as $R\perp A$, implying that the sensitive attribute has no influence on the model’s output.
    • Separation — allowing for some correlation between $A$ and the target variable $Y$ if it is justifiable, denoted as $R\perp A\mid Y$.
    • Sufficiency — ensuring that the regression output adequately captures the influence of $A$ on $Y$, aiming for parity among individuals based on their relevant characteristics, denoted as $Y\perp A\mid R$ [66].
    Additionally, fairness can be quantified using statistical measures such as the Kullback-Leibler divergence to compare predictions across different groups [30].
Each of the above metrics addresses different aspects of fairness and is applicable in various scenarios within the recruitment process. Their effective implementation can help mitigate biases and ensure a more equitable hiring process.
IV-B Bias Mitigation
After detecting bias within an AI system, the next step involves implementing strategies to mitigate these biases effectively. This subsection outlines three primary categories of bias mitigation techniques as identified in existing literature [132]:
@@ IV-B Bias Mitigation
    Pre-processing: This approach involves modifying the dataset before the model training process begins. Techniques in this category, such as reweighting and optimized pre-processing, adjust features and labels in the dataset to align with fairness criteria before the model is trained and used for prediction [132]. These methods aim to remove or modify bias-inducing factors in the data, such as unrelated protected attributes. Toolkits like AI Fairness 360 offer a range of pre-built functions for these pre-processing tasks [132]. However, while effective at data-level adjustments, pre-processing does not rectify biases inherent to the core modeling process, which may still require attention if biases persist after dataset adjustments.
@@
    Post-processing/transparency: Post-processing techniques focus on adjusting the outputs of an already-trained model to satisfy fairness criteria. For instance, this might involve modifying the classification thresholds used to make hiring decisions. The primary advantage of this approach is that it does not necessitate changes to the underlying model architecture. Additionally, these methods often incorporate elements of algorithmic transparency, such as providing explanations for decisions through counterfactual reasoning or analysis of feature importance. Revealing the decision-making process enables a thorough assessment of its fairness and allows for necessary adjustments if biases are identified. Moreover, providing clear explanations helps job candidates understand potential discrepancies or deficiencies in their applications, which can be beneficial for their job-seeking future endeavors. Research has shown that transparent feedback can significantly improve candidates’ perceptions of fairness within the recruitment process [135].
Numerous ML toolkits have been developed to facilitate the integration of fairness algorithms at various stages of the ML pipeline. These toolkits, comprehensively reviewed in the survey by Lee et al. [136], support the implementation of pre-processing, in-processing, and post-processing methods. Their availability enables organizations to more easily embed fairness into their AI-driven hiring processes, thereby helping ensure that their recruitment practices meet established fairness standards.
IV-C Observations, Limitations, and Future Challenges
Here, we outline four key limitations in AI-driven hiring that underscore the need for ongoing research to enhance the fairness of AI in recruitment processes:
@@ IV-C Observations, Limitations, and Future Challenges
    Selecting metrics and auditing practices: As organizations implement fairness practices, a pivotal challenge arises in selecting appropriate fairness metrics, which can vary significantly across different applications. In recruitment, for example, it is crucial to consider both candidate qualifications and the representation of protected groups. Additionally, the difficulty of meeting multiple fairness criteria simultaneously [134] suggests that relying on a single metric might ignore other significant disparities. Equally concerning is the absence of standardized benchmarking and comprehensive auditing procedures in current legislation, which are vital for consistently ensuring fairness in the use of AI tools in hiring [109].
@@
    Job-specific requirements and long-term fairness: The criteria for measuring fairness must adapt to the unique requirements of different jobs. For example, while religion is generally a protected attribute, it may be relevant for leadership roles within religious institutions. This underscores the challenge of using a single model to predict protected attributes universally across all job types, especially where specific attributes are crucial. Additionally, the need for fairness extends to job advertisements and their recommendations, which must align with the detailed requirements specified in the job descriptions. As the job market evolves, the models and auditing mechanisms must also be dynamic, adapting to long-term changes in the market. This dynamic aspect of fairness, focusing on long-term fairness impacts and adaptability, has been underexplored in the fairness literature, which has traditionally focused on static fairness assessments.
@@
    Auditing large foundation models: Large foundation models like ChatGPT, Whisper, and DALL-E are increasingly employed across various hiring tasks due to their robust performance in diverse scenarios. These models utilize extensive training datasets comprising text, images, and audio. However, their widespread use raises significant ethical concerns due to the current lack of comprehensive regulations that ensure fairness throughout the model training and deployment phases. Key issues include preventing the leakage of sensitive data when generative models create content and promoting diversity in training datasets to minimize the perpetuation of stereotypes and harmful biases, thereby safeguarding equal treatment of all individuals [137].
@@
    Hiring decision transparency: Many applicants never receive feedback if screened out by an AI-based system, and those who do often get responses lacking detailed justification. This lack of detailed feedback on unmet requirements or the features influencing the AI’s decision undermines interactional justice and erodes trust. Enhancing the interpretability and explainability of ’black-box’ AI models is crucial for fostering transparency and fairness in AI-driven hiring. Despite emerging legislation aimed at improving transparency in AI-driven decisions, such initiatives are still limited and often lack robust enforcement mechanisms.
These outlined challenges highlight the complex, evolving landscape of AI-driven hiring, underscoring the critical need for continued innovation and rigorous evaluation to advance fairness and transparency in recruitment practices.
V Conclusion
In this paper, we have systematically reviewed the existing literature on fairness in AI-based recruitment systems, elucidating the complexities and challenges at the intersection of AI research and development, I-O psychology, and legislative frameworks. We began by defining essential concepts such as fairness, trust, and justice, grounding our discussion in relevant literature to provide a foundational understanding for analyzing bias. We then detailed the recruitment pipeline, illustrating the application of AI at each stage to both detect and mitigate bias, and highlighted the current methodologies and their limitations as outlined in the fairness literature, specifically tailored to AI-driven recruitment.
We identified significant areas in need of future research, including the development of models that cater to job-specific requirements, the creation of fair job advertisements to engage a broader audience, and enhancing transparency in recruitment decisions. Additionally, pinpointing the origins of bias remains a pivotal challenge for the advancement of fair AI practices in recruitment.

### v3

@@ 4 Bias Detection and Mitigation
Ensuring fairness in AI-driven hiring requires systematic methods for detecting, measuring, and mitigating bias at each stage of the recruitment pipeline. In this section, we review the existing work in AI fairness, focusing on methods for measuring bias, strategies for mitigating bias, and the current limitations and challenges in this field.
4.1 Fairness Metrics
Different fairness metrics have been proposed within AI fairness research, each capturing distinct dimensions of fairness. As discussed later in this subsection, fairness metrics often involve inherent trade-offs that require careful consideration. This subsection reviews core fairness metrics that are particularly relevant for recruitment and hiring contexts, drawing upon foundational literature in algorithmic fairness [83, 90, 93].
@@ 4.1 Fairness Metrics
    Fairness through unawareness: This metric provides a baseline definition of fairness, asserting that an algorithm is fair if it does not explicitly consider protected attributes ($A$) during training or decision-making [178]. Formally, a classifier $C:X\setminus A\to Y$ is trained on feature set $X$ excluding the protected attributes $A$, where $Y$ represents possible outcomes. Although intended to avoid direct discrimination by ensuring the model is “unaware” of protected characteristics, this approach has several limitations. Most notably, features excluded from $X$ can often be indirectly inferred through other correlated attributes, termed proxies. In a hiring scenario, even if gender is not directly included, an indicator such as attendance at a women-focused coding bootcamp might implicitly reveal the applicant’s gender, thereby introducing unintended biases. Similarly, the prestige or location of educational institutions might indirectly encode socioeconomic status or ethnicity. Therefore, achieving fairness through unawareness alone is insufficient, and expert knowledge remains crucial for accurately identifying and addressing such hidden biases across different occupational contexts.
@@
    Demographic parity: Also known as statistical parity, this metric stipulates that the acceptance rates for different demographic groups should be approximately equal. This requirement aligns closely with the concept of adverse impact as defined by the EEOC’s four-fifths rule, detailed in Section 2.2. Mathematically, let $C$ be a classifier predicting a binary outcome $Y$. For any two groups $a$ and $b$, the acceptance rates must be within a specified percentage $p$ of each other, where $\epsilon=\frac{p}{100}$ and $\epsilon\in[0,1]$ [90]:
@@
    This metric emphasizes fairness by ensuring qualified candidates from all groups have an equal probability of selection. However, it does not directly address potential disparities in false-positive or false-negative rates, which can still lead to unequal outcomes across groups.
@@
    These conditions help ensure that the model’s predictions consistently reflect actual candidate qualifications across different groups, preventing systematic biases that could unfairly advantage or disadvantage specific demographics.
@@
    Counterfactual fairness: Counterfactual fairness requires that algorithmic decisions remain consistent even if a protected attribute $A$ were hypothetically changed, thus ensuring robustness against discrimination based solely on protected characteristics. Defined by Kusner et al. [178], a classifier $C$ achieves counterfactual fairness if, for any individual with features $X$ including a protected attribute $A$, the outcome remains the same in both the actual and a counterfactual scenario where $A$ is altered. The mathematical definition is:
@@
    Individual fairness (versus group fairness): Unlike group fairness metrics such as demographic parity, accuracy parity, and predictive rate parity which ensure fairness by equalizing statistical measures across groups, individual fairness focuses on fairness at the individual level. This metric asserts that similar individuals should receive similar outcomes. It addresses the key limitation of group fairness metrics, which ensure fairness across groups but may still permit unfairness within groups [179]. Individual fairness requires that any two individuals who are comparable regarding relevant qualifications and characteristics for a job should receive similar outcomes from the model [180].
@@
    Preference-informed individual fairness: This fairness metric extends individual fairness by considering the diverse preferences of individuals, such as desired job location, industry, salary expectations, and work hours. Preference-informed individual fairness (PIIF) incorporates the concept of envy-freeness, which means no individual would prefer another’s outcome over their own under similar circumstances [180]. PIIF thus requires that outcomes not only meet individual fairness standards but also respect the personal preferences of individuals, ensuring decisions made by the model are both fair and aligned with individual job-related priorities.
@@
    Multi-sided fairness: Modern recruitment platforms involve multiple stakeholders—employers posting job advertisements (producers), jobseekers (consumers), and the platform itself (the intermediary system). A multi-sided fairness approach considers fairness simultaneously from the perspective of all these groups [181]. Focusing exclusively on one stakeholder can unintentionally disadvantage others. For instance, a consumer-focused fairness approach (C-fairness) might mandate that job advertisements display similar salary ranges to all jobseekers, regardless of their protected group membership. Conversely, a producer-focused fairness approach (P-fairness) might require that advertisements from minority-owned businesses are shown to qualified jobseekers at comparable rates to those from other businesses [182, 183]. Multi-sided fairness is thus often formulated as a multi-objective optimization challenge, balancing accuracy and fairness across stakeholders, as seen in various online platforms such as e-commerce and ride-hailing services [181, 184, 182, 183].
@@
    Procedural fairness: Procedural fairness, introduced in Section 2.1, pertains to perceptions regarding the fairness of the processes used to arrive at decisions, such as candidate selection methods in recruitment. In the context of AI-based hiring, procedural fairness can be evaluated by assessing stakeholder perceptions about the legitimacy and appropriateness of features and criteria used by AI systems in making decisions. For instance, Grgic-Hlaca et al. [58] propose that procedural fairness in algorithmic decisions should reflect whether stakeholders view the use of specific features as justified and unbiased. However, perceptions of fairness can evolve as stakeholders become more aware of potential biases and unintended consequences associated with certain decision-making processes or features used by AI models.
@@
    Fairness in regression: Fairness metrics have traditionally focused on binary classification outcomes (e.g., hire or reject), yet regression models are prevalent in hiring contexts, such as predicting job performance scores from candidate interviews or resumes [185]. In fairness analyses, the actual outcome ($Y$) differs from the model’s predicted outcome ($R$), with the prediction ideally closely approximating the true outcome. Here, the protected attribute ($A$)—such as race, gender, or age—is considered when assessing potential biases in these predictions. According to Barocas et al. [86], fairness in regression is guided by three principles:
    • Independence—ensuring the protected attribute $A$ and the regression outcome $R$ are statistically independent ($R\perp A$), indicating no influence of protected attributes on predictions.
    • Separation—permitting justified correlations between the protected attribute $A$ and the target variable $Y$, but ensuring predictions $R$ remain independent of $A$ given $Y$ ($R\perp A\mid Y$).
    • Sufficiency—ensuring predictions fully represent relevant characteristics, such that outcomes $Y$ are independent of protected attributes $A$ given the predictions $R$ ($Y\perp A\mid R$) [83].
    Moreover, regression fairness can be quantitatively evaluated using metrics like the Kullback-Leibler divergence, assessing differences in predictive distributions across demographic groups [34].
Metric Selection: Each of the above metrics addresses different aspects of fairness and is applicable in various scenarios within the recruitment process to ensure an equitable hiring process. Selecting the right fairness metrics for AI-driven hiring requires careful evaluation of both organizational objectives and the inherent trade-offs between different definitions. For instance, counterfactual fairness states an algorithm is fair if the hiring decision would remain unchanged for a candidate solely if their protected attribute were altered, while demographic parity stipulates the acceptance rates for different demographic groups should be approximately equal. These metrics may compete with one another and optimizing for one can compromise another aspect of fairness or accuracy. Moreover, fairness metrics developed in academic and technical contexts do not always align perfectly with legislative requirements. Laws such as Title VII of the Civil Rights Act establish broad non-discrimination baselines without prescribing specific numerical thresholds or computational formulas, leaving room for interpretation. Therefore, it is essential to align fairness metrics with both legal standards and operational goals. A multidisciplinary approach also involving legal experts and ethicists can help organizations navigate these differences and choose or develop measures that balance multiple fairness dimensions while complying with regulatory obligations. Experts in hiring and the specific job area are also needed to identify protected attributes and potential proxies that could contribute to bias. Once these protected attributes are determined, fairness metrics should be chosen to align with the hiring task. For example, during resume screening, metrics such as demographic parity and counterfactual fairness can be used to ensure selection rates are consistent across groups; however, for video interviews or later stages, a more fine-grained metric—such as accuracy parity—may be better suited to ensure candidates are adequately qualified. Additionally, ongoing bias evaluations should be conducted, as definitions of fairness and legislation evolve over time.
4.2 Bias Mitigation Strategies
After detecting bias within an AI system, the next step involves implementing strategies for mitigation. This subsection outlines three primary categories of bias mitigation techniques as identified in existing literature [186]:
@@ 4.2 Bias Mitigation Strategies
    Pre-processing: This approach involves modifying the dataset before the model training process begins. Techniques in this category, such as reweighting and optimized pre-processing [187], adjust features and labels in the dataset to align with fairness criteria before the model is trained and used for prediction [186]. These methods aim to remove or modify bias-inducing factors in the data, such as unrelated protected attributes, potential proxies of protected attributes, or imbalances in groups between training and test splits. Toolkits like AI Fairness 360 offer a range of pre-built functions for these pre-processing tasks [188]. Other recent techniques such as generation of synthetic data to balance representation in training datasets have also been proposed [189]. However, while effective at data-level adjustments, pre-processing does not rectify biases inherent to the core modeling process, which may still require attention if biases persist after dataset adjustments.
@@
    Post-processing/transparency: Post-processing techniques focus on adjusting the outputs of an already-trained model to satisfy fairness criteria. For instance, this might involve modifying the classification thresholds used to make hiring decisions, or even modifying the predicted outcomes themselves [186]. The primary advantage of this approach is that it does not necessitate changes to the underlying model architecture. Additionally, these methods often incorporate elements of algorithmic transparency, such as providing explanations for decisions through counterfactual reasoning or analysis of feature importance. Revealing the decision-making process enables a thorough assessment of its fairness and allows for necessary adjustments if biases are identified. Moreover, providing clear explanations helps job candidates understand potential discrepancies or deficiencies in their applications, which can be beneficial for their future job searches. Research has shown that transparent feedback can significantly improve candidates’ perceptions of fairness within the recruitment process [192].
Selecting a Bias Mitigation Strategy: Upon discovering bias or unfairness in an AI system at any stage of the recruitment process, a bias mitigation strategy should be developed. The first step is to assess the underlying cause of the bias, which will guide the application of mitigation techniques. For instance, if an imbalance among groups in the dataset is contributing to bias, pre-processing techniques may be employed. However, it is important to consider the feasibility of different approaches for various researchers and practitioners. For example, pre-processing requires retraining the model, which incurs computational costs, while in-processing may not be feasible if the AI model is outsourced.
Another crucial aspect of bias mitigation is the continuous evaluation of bias even after a mitigation strategy has been applied. This ensures consistent fairness as hiring models evolve, practices change over time, and local legislation is followed. Numerous ML toolkits have been developed to integrate fairness algorithms at various stages of the ML pipeline. These toolkits—comprehensively reviewed in the survey by Lee et al. [193]—support the implementation of pre-processing, in-processing, and post-processing methods. Their availability enables organizations to more easily embed fairness into their AI-driven hiring processes, thereby helping ensure that their recruitment practices meet established fairness standards. For a more detailed overview of bias mitigation strategies, we point readers to the survey by Hort et al. [186].
5 Limitations, Challenges, and Future Directions
This section outlines key limitations and challenges in current approaches to bias detection and mitigation in AI-driven recruitment, highlighting the need for continued research and development in this domain:
@@ 5 Limitations, Challenges, and Future Directions
    Selecting metrics and auditing practices: As organizations implement fairness practices, a pivotal challenge arises in selecting appropriate fairness metrics, which can vary significantly across different applications. In recruitment, for example, it is crucial to consider both candidate qualifications and the representation of protected groups. Additionally, the difficulty of meeting multiple fairness criteria simultaneously [191] suggests that relying on a single metric might ignore other significant disparities. Equally concerning is the absence of standardized benchmarks and comprehensive auditing guidelines in existing legislation, both essential for consistently assessing and ensuring fairness in AI recruitment tools [126].
@@
    Job-specific requirements and long-term fairness: The criteria for measuring fairness must adapt to the unique requirements of different jobs. For instance, although religion is generally considered a protected attribute, it may be legitimately relevant for leadership roles within religious organizations. This underscores the complexity of employing universal AI models across diverse job types, particularly when certain protected attributes can be legitimately job-relevant. Additionally, the need for fairness extends to job advertisements and their recommendations, which must align with the detailed requirements specified in the job descriptions. As the job market evolves, the models and auditing mechanisms must also be dynamic, adapting to long-term changes in the market. This dynamic aspect of fairness, focusing on long-term fairness impacts and adaptability, has been underexplored in the fairness literature, which has traditionally focused on static fairness assessments.
@@
    Hiring decision transparency: Many applicants who are screened out by AI-based systems never receive feedback, and even when feedback is provided, it often lacks sufficient detail or clarity regarding the reasons behind the decision. This lack of detailed feedback on unmet requirements or the features influencing the AI’s decision undermines interactional justice and erodes trust. Enhancing the interpretability and explainability of black-box AI models is crucial for fostering transparency and fairness in AI-driven hiring. Despite emerging legislation aimed at improving transparency in AI-driven decisions, such initiatives are still limited and often lack robust enforcement mechanisms.
These outlined challenges highlight the complex, evolving landscape of AI-driven hiring, underscoring the critical need for continued innovation and rigorous evaluation to advance fairness and transparency in recruitment practices. To address these challenges, we propose four potential future directions:
  1. 1.
    Standards for AI Hiring: Establishing rigorous standards or legislation for AI-driven hiring in government is essential to ensure that recruitment processes are fair. As organizations increasingly turn to AI to streamline candidate screening and decision-making, these systems must be held to robust ethical and legal benchmarks with clear guidelines outlining requirements for candidate rights to data privacy, using transparent AI-driven decisions with explanations, and implementing bias mitigation techniques. The EU’s proposed AI Act [67] already categorizes AI usage in hiring as a high‐risk application, highlighting the need for rigorous regulatory standards. Implementing similar legal frameworks in other jurisdictions is essential for establishing globally consistent standards for fair AI-driven recruitment. The Algorithmic Accountability Act of 2022, that requires companies to assess the impact of their automated decision-making systems, is a step forward, although is still in its proposal stages and does set forward explicit guidelines for nuanced bias measurement and mitigation in hiring [76].
  2. 2.
    Bias auditing large pre-trained models: Large foundation models like ChatGPT, Whisper, and DALL-E are increasingly employed across various hiring tasks due to their robust performance in diverse scenarios. These models utilize extensive training datasets comprising text, images, audio, and video data, each modality introducing distinct fairness challenges and biases. However, their widespread use raises significant ethical concerns due to the current lack of comprehensive regulations that ensure fairness throughout the model training and deployment phases. Key issues include preventing the leakage of sensitive data when generative models create content and promoting diversity in training datasets to minimize the perpetuation of stereotypes and harmful biases, thereby safeguarding equal treatment of all individuals [194].
  3. 3.
    Extend access to AI: Extending access, or democratizing, AI tools to all researchers—even those with limited resources—is crucial for improving fairness in AI-driven hiring practices. When a diverse pool of researchers can independently audit, test, and refine AI hiring systems, it ensures that bias mitigation techniques are reflective of a wide array of perspectives [195]. However, to train state-of-the-art large foundation models, there is a large and growing computational, financial, and environmental cost barrier to entry for many researchers [195]. Open access, or open-sourcing of models, accompanied by clear ethical usage guidelines and transparency requirements, can foster broader community participation, enabling independent audits that identify and mitigate discriminatory practices embedded within proprietary AI systems.
@@
    Including human feedback with AI: Ensuring that human judgment remains a vital part of the hiring process is crucial to prevent the pitfalls of fully automated, AI-driven systems. By mandating that a certain percentage of hiring decisions involve human oversight, such as human resource experts, auditors, and hiring managers, organizations can mitigate potential biases in algorithmic models and offer nuanced decisions that models may miss. Properly trained human evaluators can provide contextual insights and ethical judgment that AI systems often struggle to capture fully, thereby facilitating a more holistic and fair assessment of candidates. This collaborative approach creates a continuous feedback loop, enabling AI systems and human evaluators to iteratively refine their practices and adapt to evolving standards of fairness and ethical hiring.
6 Conclusion
In this paper, we systematically reviewed existing research on algorithmic fairness in AI-based recruitment systems, highlighting key complexities and challenges at the intersection of artificial intelligence, industrial-organizational psychology, and employment law. We began by clarifying essential concepts, including fairness, trust, and justice, grounding our discussion in interdisciplinary literature to establish a robust foundation for analyzing bias in AI hiring. We then examined the recruitment pipeline, discussing the role of AI at each stage, methods for detecting and mitigating bias, and the existing limitations and challenges highlighted in recent fairness literature.
We identified significant areas in need of future research, including the development of AI models that incorporate fairness metrics tailored to job-specific requirements, the creation of fair job advertisements to engage a broader audience, and enhancing transparency in recruitment decisions. Additionally, pinpointing the origins of bias remains a pivotal challenge for the advancement of fair AI practices in recruitment.